OpenStudy (anonymous):

a 5.05 gram sample of a copper substance was heated in air and 6.32g of copper II oxide was formed. What was the formula of the original copper substance?

OpenStudy (anonymous):

We're not sure if the copper is pure so we use the following: \[2CuX+O_2 => 2CuO +2X\] To determine the amount of O_2 added we subtract the weight of the copper substance from the weight of the Copper II Oxide. \[6.32g-5.05g=1.27g O_2\] Now we can determine the moles of O_2: \[1.27gO_2\frac{1 mol O_2}{31.9988g O_2}=0.0397molO_2\] Now that we have the moles of O_2 we can calculate the moles of Cu. \[0.0794molCu\frac{63.546gCu}{1molCu}=5.05gCu\] Not this is the value of our original substance so the original substance is pure copper.

OpenStudy (anonymous):

*Note